# When Respondents Don’t Care Anymore: Identifying the Onset of Careless Responding

### v2

@@ B.3 Details on Changepoint Detection
Appendix C Details on the Empirical Application
C.1 Data Preprocessing
The data of Johnson, (2005), provided at https://osf.io/sxeq5/, contain the following information for each of the the $n=20,993$ respondents. First, per-item responses to $p=300$ IPIP-NEO-300 items, five domain scores (directly calculated from the responses), scores for psychometric antonyms and personal reliability, as well as respondent sex and age.
The responses in the provided data have already underwent some preprocessing that is described in detail the accompanying README file. Specifically, responses to negatively-worded items were reverse-coded, that is, a given response of 5 was recoded to 1, 4 was recoded to 3, and so on. Hence, in order to recover the original responses, we reverted this recoding. Furthermore, a total of 33,691 individual responses are missing (0.53% of all responses). Note that Johnson, had already removed participants with too many missing responses from the provided data set, hence there are relatively few missing responses for the remaining participants. Due to the low number of missing responses, we treat them as a separate answer category by coding them as 0.
C.2 Additional Results
Figure C.1 visualizes the detected onset items for the respondents that are flagged as careless by CODERS at significance levels 0.1%, 0.5%, and 1%.
Figure C.1: Estimated onset of careless responding for participants flagged by CODERS in the data of Johnson, (2005).
Note. Results for different significance levels $\alpha\in\{0.01,0.05,0.1\}$ are presented in separate panels. The $x$-axis shows the indices of the participants flagged as partially carelessness by CODERS. The participants are sorted in ascending order of the estimated carelessness onset item, which is in turn depicted on the $y$-axis. The dashed horizontal line represents the average location of carelessness onset across the flagged participants.
Figure C.2 provides a histogram as well as a density estimate for the per-respondent average reconstruction errors. We can see that the density has a long right tail, which might be due to careless respondents. We highlight the value 0.064 by a vertical dotted line, which is the average reconstruction error of the participant from Figure 3(d) and Table 2(d). This value is in the 99.7% percentile of all $n=20,993$ average reconstruction errors. Given the extremely high average reconstruction error and the fact that this respondent was flagged by both benchmark methods, it is likely that the participant responded carelessly throughout the survey, explaining why CODERS did not flag a changepoint.
Figure C.2: Histogram of average reconstruction errors for the data of Johnson, (2005).
Note. The $x$-axis contains the per-respondent average reconstruction errors, $\textsf{RE}_{i}=\frac{1}{p}\sum_{j=1}^{p}\textsf{RE}_{ij}$, $i=1,\dots,n$, where averages are taken over the $p=300$ items. On the $y$-axis, a histogram of the average reconstruction errors for $n=20,993$ respondents is displayed together with an empirical Gaussian kernel density. The vertical dotted line denotes the value of $\textsf{RE}_{i}$ for the participant in Figure 3(d) that was flagged by both benchmark methods, but not CODERS. This value, 0.064, lies in the 99.7% percentile of all average reconstruction errors.
Appendix D Details and Additional Simulations
D.1 Details on the Adjusted Rand Index
In the context of our simulations, the adjusted Rand index (ARI) for a given respondent is computed as
$ARI=\begin{cases}1&\text{ if the respondent is attentive and no changepoint isflagged},\\ \frac{A-BC/D}{(B+C)/2-BC/D}&\textnormal{ otherwise,}\end{cases}$
where
$A$ $=\binom{p_{11}}{2}+\binom{p_{00}}{2}+\binom{p_{10}}{2}+\binom{p_{01}}{2},$
$B$ $=\binom{p_{00}+p_{01}}{2}+\binom{p_{10}+p_{11}}{2},$
$C$ $=\binom{p_{00}+p_{10}}{2}+\binom{p_{01}+p_{11}}{2},$
$D$ $=\binom{p}{2},$
with $p_{11}$ denoting the number of careless responses that are correctly identified as careless, $p_{00}$ the number of attentive responses that are correctly identified as attentive, $p_{10}$ the number of careless responses that are incorrectly identified as attentive, and $p_{01}$ the number of attentive responses that are incorrectly identified as careless. Intuitively, $p_{11}$ and $p_{00}$ measure correct classification, whereas $p_{01}$ and $p_{10}$ measure misclassification. Observe that the ARI assumes value 1 if there is no misclassification, which is the maximum value it can take. Hence, for accurate estimates of the carelessness onset item, the ARI will be close to value 1. For very inaccurate estimates of carelessness onset, the ARI will be close to 0, or potentially even below that. It is unclear how to interpret negative ARI values, but such values only occur rarely. We refer to Steinley, (2004) for a detailed discussion of the properties of the ARI.
If a changepoint is flagged in an attentive respondent, the ARI takes value 0. Similarly, if no changepoint is flagged in an attentive respondent, the ARI takes value 1. Hence, when averaging the ARI over attentive respondents, the ensuing average can be interpreted as the proportion of attentive respondents who are correctly identified as attentive. On the other hand, averages of ARIs calculated on partially careless respondents do not have a direct interpretation other than higher values indicating better performance.
D.2 Additional Simulation Results
We investigate the effect of various choices in our simulations. As the results are similar to those in the main paper other than some shifts in size or power, we only present plots of the average ARI for attentive and careless respondents (similar to Figure 5) in order to keep the appendix concise.
D.2.1 Choosing a Different Significance Level
Figure D.1: ARI of CODERS at a significance level of 0.5%.
Note. The ARI is averaged over fully attentive respondents and partially careless respondents (columns) for different carelessness onset settings (rows).
Figure D.2: ARI of CODERS at a significance level of 1%.
Note. The ARI is averaged over fully attentive respondents and partially careless respondents (columns) for different carelessness onset settings (rows).
We use the baseline simulation design from the main text, but we set the significance level in CODERS to 0.5% and to 1%, respectively. The corresponding results are shown in Figures D.1 and D.2. Compared to the significance level of 0.1% in the main text (cf. Figure 5), there is a gradual increase in power in the changepoint detection, yielding a small increase in the ARI across prevalence levels for careless respondents. Further investigation suggests that the increase in ARI stems mainly from better performance in detecting random and extreme responding (results not shown to keep the appendix at a reasonable length). However, for high prevalence levels from 40% onwards, this comes at a cost of too many false positives in attentive respondents, translating to a substantial decrease in ARI for the highest prevalence levels.
D.2.2 Varying Correlation Structures
In this variation of the simulation design, we keep the same settings as in the baseline design from the main text, but we vary the correlation structure in the following ways:
  1. 1.
    Randomly sampled moderate correlations. In every repetition, correlations of item pairs from the same scale are randomly drawn from the interval $[0.4,0.6]$. Items from different scales remain uncorrelated.
  2. 2.
    Randomly sampled strong correlations. In every repetition, correlations of item pairs from the same scale are randomly drawn from the interval $[0.6,0.8]$. Items from different scales remain uncorrelated.
  3. 3.
    Randomly sampled grouped correlation structure. Multiple scales form groups so that there are 5 groups consisting of 6 scales each. Since there are 10 items in each scale, there are $6\times 10=60$ items in each group. In each repetition, correlations of item pairs from the same scale are randomly drawn from the interval $[0.6,0.8]$, correlations of item pairs in different scales of the same group are randomly drawn from the interval $[0,0.2]$, and items from different groups are uncorrelated.
While the baseline simulation design uses a relatively simple correlation structure (items from the same scale have a fixed correlation of 0.7 before recoding, items from different scales are uncorrelated), the three variations above introduce more randomness into the design. Due to this randomness, it may occur that a sampled correlation matrix is not positive definite. In such a case, we find the closest positive definite correlation matrix via the function nearPD() in the R package Matrix (Bates et al.,, 2023).
Figure D.3: ARI of CODERS for randomly sampled moderate correlations.
Note. The ARI is averaged over fully attentive respondents and partially careless respondents (columns) for different carelessness onset settings (rows).
Figure D.4: ARI of CODERS for randomly sampled strong correlations.
Note. The ARI is averaged over fully attentive respondents and partially careless respondents (columns) for different carelessness onset settings (rows).
Figure D.5: ARI of CODERS for the randomly sampled grouped correlation structure.
Note. The ARI is averaged over fully attentive respondents and partially careless respondents (columns) for different carelessness onset settings (rows).
Figures D.3D.5 display the results corresponding to the different correlation structures at significance level 0.1%. The results of the design with randomly sampled strong correlations (Figure D.4) and the grouped correlation structure (Figure D.5) are essentially equivalent to the results of the baseline design with fixed correlation (Figure 5 in the main text). The design with randomly sampled moderate correlations (Figure D.3) also yields similar results, save for a small power drop in the changepoint detection (resulting in a lower ARI for careless respondents) and fewer false positives in extreme carelessness prevalence levels (yielding a higher ARI in attentive respondents).
D.2.3 Varying Sample Size
This variation of the simulation design again uses the same settings as in the baseline design from the main text, but now we change the sample size to $n=250$ and $n=1000$, respectively.
Figures D.6 and D.7 visualize the corresponding results at significance level 0.1%. They are similar to the results for $n=500$ in the main text (cf. Figure 5), other than a small decrease ARI in careless respondents for $n=250$ as well as a small increase for $n=1000$.
Figure D.6: ARI of CODERS for sample size $n=250$.
Note. The ARI is averaged over fully attentive respondents and partially careless respondents (columns) for different carelessness onset settings (rows).
Figure D.7: ARI of CODERS for sample size $n=1000$.
Note. The ARI is averaged over fully attentive respondents and partially careless respondents (columns) for different carelessness onset settings (rows).

### v3

@@ B.4 Details on Changepoint Detection
Appendix C Additional Simulations
C.1 Additional Results for the Main Design
In addition to CODERS, we apply three traditional methods for the detection of careless respondents in the simulation design from the main text, which is inspired by the data collected by Johnson, (2005): personal reliability, psychometric antonym, and the longstring index. We choose these three methods since they were also applied in the study of Johnson, (2005). Following Johnson,, we reverse psychometric antonym scores, and we flag respondents as careless whose personal reliability and psychometric antonym scores are lower than $0.3$ and $-0.03$, respectively. For the longstring index, we flag a respondent as careless if its value exceeds 6 (i.e., $>6$), which is a cutoff value suggested by Niessen et al., (2016).
Figure C.1 visualizes the false positive rate (FPR) and the false negative rate (FNR) of the three methods for the different carelessness onset regimes (in separate columns). Figure C.2 further shows the false negative rates for the four types of partially careless respondents (in separate rows). The longstring index performs perfectly at its intended task of detecting straightlining, whereas it hardly detects any random responding or pattern responding (which is expected). Furthermore, it performs better than expected at identifying extreme responding. Since the survey is lengthy and those respondents choose randomly from only the two extreme answer categories, there is a relatively high probability (depending on the onset) that they select a long string of the same answer category somewhere by chance. It turns out that personal reliability flags hardly any respondents, resulting in false positive rates close to 0 and false negative rates close to 1. Psychometric antonym yields inflated false positive rates, while the false negative rates are still relatively high. Interestingly, its performance is fairly constant across the different types of carelessness. As expected, the false negative rate of the traditional methods in general worsens the later the onset of careless responding (while that of CODERS deteriorates the earlier the onset, see Figures 7 and 8 in the main text).
Figure C.1: Performance measures for the traditional methods used by Johnson, (2005).
Note. The $x$-axis displays various prevalence levels. False positive rates in attentive respondents (top row) and false negative rates in partially careless respondents (bottom row) for different onset regimes (columns) are averaged across 100 repetitions.
Figure C.2: False negative rate of the traditional methods used by Johnson, (2005) for different types of partially careless respondents.
Note. The $x$-axis displays various prevalence levels. Results for different types of careless responding (rows) in different onset regimes (columns) are averaged across 100 repetitions.
C.2 Decreasing Survey Length
Recall that the changepoint detection method of Shao & Zhang, (2010) used in CODERS is based on asymptotic arguments for the number of items being large (see Appendix B.4). It therefore seems plausible that the ability of CODERS to accurately identify partial carelessness may decline as survey length decreases.
C.2.1 Data Generation
We use the same simulation design of the main text, but we focus on the baseline onset regime and instead decrease the number of items $p$ by gradually removing sets of items. Specifically, for a given simulated dataset, we remove all items belonging to randomly selected personality traits. There is a total of five traits, each of which is measured by six facets comprising ten items each (300 items in total). Hence, removing a single trait from the measurement will remove $6\times 10=60$ items. We drop up to four traits, resulting in surveys comprising a total of $p\in\{240,180,120,60\}$ items.
C.2.2 Results
Figure C.3 plots the false positive rates for the different survey lengths (in separate columns). As before, CODERS rarely incorrectly flags attentive respondents as careless at any of the considered significance levels, although there is some overrejection for the shortest survey with $p=60$. Figure C.4 visualizes the false negative rate (top row) and the mean absolute error of onset estimation (bottom row) for careless respondents across the different surveys lengths (in separate columns). At a given significance level, the false negative rate slightly increases with decreasing survey length compared to the results from the main text. For $p=60$, the false negative rate reaches about 30%–40%, depending on the significance level. In addition, the location of the onset of carelessness is still estimated highly accurately throughout the investigated survey lengths, on average being off by only about one or two items.
Figure C.3: False positive rate of CODERS for attentive respondents across decreasing survey length.
Note. The $x$-axis displays various prevalence levels of partially careless respondents. Results for different survey lengths are shown in separate columns, averaged across 100 repetitions.
Figure C.4: Performance measures of CODERS for partially careless respondents across decreasing survey length.
Note. The $x$-axis displays various prevalence levels. The top row shows results for the false negative rate, while the bottom row visualizes the accuracy of onset estimation via the mean absolute error. Results for different survey lengths are shown in separate columns, averaged across 100 repetitions.
Figure C.5: False negative rate of CODERS for different types of partially careless respondents across decreasing survey length.
Note. The $x$-axis displays various prevalence levels. Results for different types of careless responding (rows) in different survey lengths (columns) are averaged across 100 repetitions.
Figure C.6: Mean absolute error of carelessness onset estimation with CODERS for different types of partially careless respondents across decreasing survey length.
Note. The $x$-axis displays various prevalence levels. Results for different types of careless responding (rows) in different survey lengths (columns) are averaged across 100 repetitions.
Figures C.5 and C.6 show the false negative rate and the mean absolute error of onset estimation, respectively, grouped by carelessness type (rows) for each considered survey length (columns). In essence, the slight increase in false negative rates seems to be primarily caused by fewer inconsistent careless respondents being detected, whereas the performance for invariable careless responding stays near-perfect. For $p=60$, in particular very few random respondents are correctly identified, but the performance is still acceptable for extreme respondents.
We conclude that, as expected, the power of CODERS for detecting partially careless respondents drops as the survey length decreases. Yet, the drop in power is small and primarily caused by a low detection performance for random respondents. It is conceivable that reconstruction errors are a noisier measurement of inconsistency (compared to LSP as a measurement of invariability), as the autoencoder has less information to learn attentive responding behavior in shorter surveys, explaining why the loss of power is primarily visible for random responding. Nevertheless, if a changepoint is flagged, its accuracy is nearly unchanged at excellent levels. That is, even though CODERS has somewhat less power in identifying partially careless respondents, the onset of carelessness in the flagged respondents is sill accurately estimated. Furthermore, the power loss can be partially alleviated by choosing a less conservative significance level, such as 0.5% or 1%.
C.3 Additional Sample Sizes
Although the number of items in the survey may be the primary driver regarding power of the changepoint detection test used in CODERS, the autoencoder may also benefit from more information in the form of a larger sample size in order to learn attentive response behavior (and conversely suffer under a smaller sample size).
C.3.1 Data Generation
[…]
C.3.2 Results
Figures C.7C.10 visualize the performance measures for $n=248$ and $n=1,000$ simulated participants (in separate columns). The results are similar to those in the main text, with the overall FNR in partially careless respondents increasing only slightly for the smaller sample size and decreasing only slightly for the larger sample size. For random respondents, however, the increase in false negatives in the smaller sample size is notable, but so is the decrease in FNR in the larger sample size (see Figure C.9, top row). This can likely be explained by the amount of information the autoencoder has available in order to learn to distinguish between attentive responding and random responding. It is therefore conceivable that the power for detecting random respondents may improve further in even larger sample sizes.
Figure C.7: False positive rate of CODERS for attentive respondents across additional sample sizes.
Note. The $x$-axis displays various prevalence levels of partially careless respondents. Results for different sample sizes are shown in separate columns, averaged across 100 repetitions.
Figure C.8: Performance measures of CODERS for partially careless respondents across additional sample sizes.
Note. The $x$-axis displays various prevalence levels. The top row shows results for the false negative rate, while the bottom row visualizes the accuracy of onset estimation via the mean absolute error. Results for different sample sizes are shown in separate columns, averaged across 100 repetitions.
Figure C.9: False negative rate of CODERS for different types of partially careless respondents across additional sample sizes.
Note. The $x$-axis displays various prevalence levels. Results for different types of careless responding (rows) in different sample sizes (columns) are averaged across 100 repetitions.
Figure C.10: Mean absolute error of carelessness onset estimation with CODERS for different types of partially careless respondents across additional sample sizes.
Note. The $x$-axis displays various prevalence levels. Results for different types of careless responding (rows) in different sample sizes (columns) are averaged across 100 repetitions.
C.4 Randomly Choosing from the Middle Answer Categories
Recall that the simulation design in the main text considers four types of careless responding: random responding, extreme responding, pattern responding, and straightlining. However, there may be other relevant types of carelessness beyond those four, including ones that are extremely difficult to distinguish from attentive response behavior.
C.4.1 Data Generation
As before, we take the simulation design from the main text with the baseline onset regime, but instead of the four aforementioned types of careless responding, we generate a distinct type that we refer to as middling. Middling is characterized by randomly choosing between the three middle answer categories (out of five). As such, middling carelessness can be seen as an inconsistent response type complementing extreme responding. The major difference to extreme carelessness is that central response options are more frequently chosen by attentive respondents in our simulation design. In fact, with the chosen marginal distributions and correlation structure (see Figure 5), there is a lot of natural variation among the central response options in attentive responses. Hence, it is difficult to systematically distinguish middling carelessness from attentive responding, but there is also the risk that attentive respondents might be mistaken for middling careless respondents.
C.4.2 Results
Figures C.11 and C.12 visualizes the performance of CODERS when partially careless respondents are of the middling type. One the one hand, CODERS still almost never incorrectly flags an attentive respondent as partially careless. On the other hand, it identifies very few middling respondents with a false negative rate between about 85% and 95%, depending on the significance level. Hence, these simulations further demonstrate that CODERS requires inconsistent responding to be sufficiently different from attentive responding in order to distinguish between the two.
Figure C.11: False positive rate of CODERS for attentive respondents in the presence of partially careless respondents that randomly choose from the middle answer categories.
Note. The $x$-axis displays various prevalence levels of partially careless respondents. Results are averaged across 100 repetitions.
Figure C.12: Performance measures of CODERS for partially careless respondents that randomly choose from the middle answer categories.
Note. The $x$-axis displays various prevalence levels. The top row shows results for the false negative rate, while the bottom row visualizes the accuracy of onset estimation via the mean absolute error. Results are averaged across 100 repetitions.
In order to further highlight the difficulty of detecting middling carelessness, we again apply the traditional methods personal reliability, psychometric antonym, and the longstring index. The false positive rates and false negative rates are shown in Figure C.13. As in the simulation design from the main text, personal reliability flags hardly any respondents, resulting in false positive rates close to 0 and false negative rates close to 1. The performance of the longstring index is similar, which is not surprising because it was designed for detecting straightlining (invariable carelessness) rather than inconsistent responding like middling. Psychometric antonym identifies more middling respondents than CODERS with a false negative rate of about 65%–75%, but it exhibits an inflated false positive rate of up to about 15%.
Figure C.13: Performance measures of the traditional methods for partially careless respondents that randomly choose from the middle answer categories.
Note. The $x$-axis displays various prevalence levels. False positive rates in attentive respondents (top row) and false negative rates in partially careless respondents (bottom row) are averaged across 100 repetitions.
C.5 Temporary Carelessness
It seems plausible that more than one changepoint may occur in a participant’s response behavior. For instance, a respondent may start the survey attentively, but become careless at some point due to being distracted, only to start responding attentively again once the distraction vanishes.
C.5.1 Data Generation
We use the same simulation design from the main text, but instead of simulating only one changepoint, we simulate temporary carelessness as follows. We randomly sample the first changepoint from the item set $\{30,31,\dots,149,150\}$ (between 10% and 50% of all items) and the second changepoint from the item set $\{151,152,\dots,269,270\}$ (between 50% and 90% of all items). Between these two changepoints, we generate careless responses using the same four types of carelessness from the main text.
C.5.2 Results
Figures C.14 and C.15 visualize the performance of CODERS and Figures C.16 and C.17 that of the traditional methods. Clearly, CODERS rarely flags any participants, with a false positive rate close to 0 and a false negative rate close to 1 (we therefore omit separate results for the four types of partially careless responding). The same behavior is observed for personal reliability. Psychometric antonym achieves a better false negative rate (about 70%–85%) but struggles with many false positives (up to about 15%). The longstring index yields somewhat better performance, with a relatively low false positive rate and a false negative rate around 65%. As in the simulation design from the main text, it successfully captures straightliners as well as some extreme respondents (see Appendix C.1 for a more detailed discussion). In summary, CODERS is simply not designed for detecting more than one changepoint in respondent behavior, such that an extension to multiple changepoint detection may be a fruitful direction for further research.
Figure C.14: False positive rate of CODERS for attentive respondents in case of temporary carelessness.
Note. The $x$-axis displays various prevalence levels of partially careless respondents. Results are averaged across 100 repetitions.
Figure C.15: Performance measures of CODERS for partially careless respondents in case of temporary carelessness.
Note. The $x$-axis displays various prevalence levels. The top row shows results for the false negative rate, while the bottom row visualizes the accuracy of onset estimation via the mean absolute error. Results are averaged across 100 repetitions.
Figure C.16: Performance measures for the traditional methods in case of temporary carelessness.
Note. The $x$-axis displays various prevalence levels. False positive rates in attentive respondents (top row) and false negative rates in partially careless respondents (bottom row) are averaged across 100 repetitions.
Figure C.17: False negative rate of the traditional methods for different types of partially careless respondents in case of temporary carelessness.
Note. The $x$-axis displays various prevalence levels. Results for different types of careless responding are shown in separate columns, averaged across 100 repetitions.
C.6 Organizational Survey
Until here, all simulations are based on a measurement of the Big Five personality traits. To study how CODERS fares in surveys of different nature, we further apply a simulation design based on a study from organizational research.
C.6.1 Data Generation
We simulate a hypothetical survey inspired by a study conducted among university students working in teams on a project for a Strategic Management course. The measurement includes various individual- and team-level constructs, among others the HEXACO-60 personality assessment of Ashton & Lee, (2009), the goal orientation scales of Vandewalle, (1997), the transactive memory systems scales of Lewis, (2003), the intra-group conflict scales of Jehn, (1995), the intra-unit justice climate scales of Li & Cropanzano, (2009), a team commitment scale based on Mowday et al., (1979), and the team performance scale of Hackman, (1986). In total, the hypothetical survey includes $p=198$ items measuring $q=34$ constructs (which we set as the number of nodes in the bottleneck layer of the autoencoder in CODERS).
Figure C.18: Correlation matrix and marginal response distributions used in the simulated organizational survey.
(a) Correlation matrix of the $p=198$ items. A more saturated blue (red) in a cell corresponds to a stronger positive (negative) correlation between the corresponding items.
(b) Marginal response distributions of the $p=198$ items ($x$-axis), each of which has five response options ($y$-axis). A darker shade of gray indicates a higher probability of the corresponding response option.
Note. The correlation matrix and the marginal response distributions are calculated from empirical data. Of the $p=198$ items, 152 are positively scored and 46 are negatively scored. The items in both plots are presented in an order such that items measuring the same construct are adjacent.
Although team-level constructs are measured, we simplify the data generating process and simulate individual-level data using the correlation matrix and marginal distributions displayed in Figure C.18, which are based on empirical data from 322 students (after removing 29 respondents that contain missing values). Unfortunately, we are not at liberty to share the original data. Of the $p=198$ items, 152 are positively scored and 46 are negatively scored. The different constructs are clearly visible along the diagonal of the correlation matrix, as are various relationships between many of the measured constructs in the off-diagonal parts (cf. also Alfons et al., 2022a, ; Alfons et al., 2022b, ). We randomize the order of the items in a way that ensures that items measuring the same construct remain adjacent, which is better suited for surveys in organizational contexts. That is, we first randomize the order of the items within each construct, and we randomize the order of (groups of related) constructs.
Otherwise, we keep the same design as in the simulations from the main text. In particular, we simulate responses for $n=500$ participants, using the same settings for generating partially careless respondents.
C.6.2 Results
Figure C.19 shows the false positive rates for attentive respondents. Compared to the simulations from the main text, false positive rates are inflated for high prevalence levels, but they remain reasonably close to the nominal level while carelessness prevalence does not exceed 40% (in the early onset regime) or 30% (for the other onset regimes). A possible explanation for this finding may be that many measured constructs do not use negatively-scored items while others do, such that response behavior may be naturally different in some parts of the survey than in others.
Figure C.19: False positive rate of CODERS for attentive respondents in the simulated organizational survey.
Note. The $x$-axis displays various prevalence levels of partially careless respondents. Results for different onset regimes are shown in separate columns, averaged across 100 repetitions.
In Figures C.20C.22, the performance measures for (different types) of partially careless respondents are displayed. The overall false negative rate is similar to that of the simulations from the main text, despite the shorter survey length. But interestingly, there are slightly fewer false negatives for random responding and slightly more false negatives for extreme responding. This may be explained by differences in response distributions, with higher skewness in some of the measurements compared to the data by Johnson, (2005) (cf. Figure C.18 with Figure 5 from the main text). In addition, the onset of careless responding is estimated with similar accuracy, with an error of about two to three items averaged across the different types of partially careless responding. All in all, these results are an indication that the performance of CODERS may not depend too much on the institutional setting of the survey.
Figure C.20: Performance measures of CODERS for partially careless respondents in the simulated organizational survey.
Note. The $x$-axis displays various prevalence levels. The top row shows results for the false negative rate, while the bottom row visualizes the accuracy of onset estimation via the mean absolute error. Results for different onset regimes are shown in separate columns, averaged across 100 repetitions.
Figure C.21: False negative rate of CODERS for different types of partially careless respondents in the simulated organizational survey.
Note. The $x$-axis displays various prevalence levels. Results for different types of careless responding (rows) in different onset regimes (columns) are averaged across 100 repetitions.
Figure C.22: Mean absolute error of carelessness onset estimation with CODERS for different types of partially careless respondents in the simulated organizational survey.
Note. The $x$-axis displays various prevalence levels. Results for different types of careless responding (rows) in different onset regimes (columns) are averaged across 100 repetitions.
Appendix D Details on the Empirical Application
D.1 Data Preprocessing
[…]
D.2 Additional Results
[…]
Figure D.1: Histogram of average reconstruction errors for the data of Johnson, (2005).
[…]